# The growth of magnetic energy during the nonlinear phase of the subsonic and supersonic small-scale dynamo

### v1

@@ Abstract
Small-scale dynamos (SSDs) amplify magnetic fields in turbulent plasmas. Theory predicts nonlinear magnetic energy growth $E_{\mathrm{mag}}\propto t^{p_{\mathrm{nl}}}$, but this scaling has not been tested across flow regimes. Using a large ensemble of SSD simulations spanning sub- to supersonic turbulence, we find linear growth ($p_{\mathrm{nl}}=1$) in subsonic flows and quadratic growth ($p_{\mathrm{nl}}=2$) in supersonic flows. In all cases, we find a dynamo efficiency of $\sim 1/100$ and a duration $\Delta t\approx 20\,t_{0}$, with $t_{0}$ the turnover time, establishing the universal timescales and efficiencies of the nonlinear SSD across astrophysical and laboratory plasmas.
@@ I Introduction
Small-scale dynamo (SSD) action describes the process by which motions in a plasma with dynamically weak magnetic fields amplify and then maintain the magnetic energy density, $E_{\mathrm{mag}}$, to levels comparable to the turbulent kinetic energy density, $E_{\mathrm{kin}}$ [1]. This process is ubiquitous across space, astrophysical, geophysical, and laboratory environments, both those where the plasma supports turbulent subsonic motions (turbulent sonic Mach number $\mathcal{M}=u_{0}/c_{s}<1$, where $u_{0}$ is the rest-frame root-mean-squared velocity and $c_{s}$ the sound speed) and where they are supersonic ($\mathcal{M}>1$). The fields generated in this process magnetize the plasma between galaxies [2, 3], provide pressure support against collapse in galaxy mergers [4], enable rapid spin-down of stellar merger remnants [5], and modify cosmic ray propagation via curvature acceleration in the interstellar medium (ISM) [6, 7]. Direct evidence for SSD action comes both from in situ observations in the Earth’s subsonic magnetosheath [8] and from supersonic laboratory laser experiments [9, 10, 11, 12].
@@
SSDs evolve through a number of distinct phases [13, 14, 15, 16, 17, 18, 1]. At first the magnetic field is too weak to exert significant forces on the plasma and thus the velocity field is kinematic, which leads to exponential amplification of the magnetic energy density, $E_{\mathrm{mag}}\propto\exp(\gamma_{\mathrm{exp}}t)$, where $t_{0}=\ell_{0}/u_{0}$ is the outer-scale eddy turnover time in a flow with outer scale $\ell_{0}$ [13, 15]. The growth rate scales are predicated to scale as $\gamma_{\mathrm{exp}}t_{0}\propto\mbox{Re}^{1/2}$ for $\mathcal{M}<1$ and $\gamma_{\mathrm{exp}}t_{0}\propto\mbox{Re}^{1/3}$ for $\mathcal{M}>1$, where $\mbox{Re}\sim u_{0}\ell_{0}/\nu$ is the hydrodynamic Reynolds number [19] and $\nu$ is the kinematic viscosity. Field amplification is driven by chaotic stretching, $u_{\ell}/\ell$ [1], which occurs on an eddy turnover timescale, $t_{\ell}$, and thus is fastest on the viscous scale $\ell_{\nu}$, which, for Kolmogorov-turbulence turbulence, is the scale on which eddies evolve on the shortest timescales [20, 21, 22, 23, 24, 25]. However, once the kinematic dynamo amplifies the field such that $E_{\mathrm{mag}}\approx E_{\mathrm{kin}}$ on scale $\ell_{\nu}$, the magnetic field begins to back react on the velocity, marking the end of the kinematic phase and the start of the nonlinear phase. During this phase the dominant stretching scale $\ell_{s}$ shifts to larger $\ell$ (longer wavelength $k$) with longer $t_{\ell}$, driving a slower, secular dynamo. As the field grows, it suppresses $u_{\ell}/\ell$ on ever-larger scales, until $\ell_{s}$ reaches a maximum scale. At this point the dynamo reaches the saturated phase, and becomes statistically steady, driven MHD turbulence [26].
@@
Both numerical simulations [17, 23, 24], and laboratory experiments [10, 11, 12] confirm that the evolution of $E_{\mathrm{mag}}$ during the kinematic phase agrees well with both the Kazanstev, Anderson & Kulsrud SSD theory [27, 15] as well as with numerical works in the $\mathcal{M}>1$ regime [28, 29, 24]. However, this well-understood phase likely ends almost immediately in real astrophysical plasmas. This is because the growth rate is $\gamma_{\mathrm{exp}}t_{0}\propto\mbox{Re}^{1/2}$, and many warm or cold astrophysical plasmas, at least in our Galaxy and the surrounding medium, often have $\mbox{Re}\approx 10^{3}-10^{10}$, thus making the kinematic phase persist for only a small fraction of the outer dynamical timescale. Hence, almost any observable astrophysical system where the field is growing will instead be in the nonlinear phase, which has not been confronted with detailed numerical measurements like the kinematic dynamo.
@@
While there is indication that this phase yields secular growth in the magnetic energy [30, 31, 32, 33, 24], $E_{\mathrm{mag}}=\alpha_{\mathrm{nl}}t^{p_{\mathrm{nl}}}$, there are no detailed calculations measuring the efficiency $\alpha_{\mathrm{nl}}$ and growth order $p_{\mathrm{nl}}$ across a broad range of plasma parameters. For incompressible Kolmogorov-like turbulence, where $E_{\mathrm{kin}}(k)\propto k^{-5/3}$, the prevailing expectation and hint from numerical simulations is that $p_{\mathrm{nl}}=1$ [15, 16, 31, 30, 34], and $\alpha_{\mathrm{nl}}\propto\varepsilon$, where $\varepsilon\sim\rho_{\ell}u_{\ell}^{3}/\ell$ is the energy flux rate directly from the hydrodynamical cascade. In contrast, for highly-compressible, shock-dominated turbulence, where $E_{\mathrm{kin}}(k)\propto k^{-2}$ as in Burgers-like turbulence [35, 36, 24, 25], the outcome is debated: some models [37, 38] predict quadratic growth ($p_{\mathrm{nl}}=2$), while others suggest universal linear growth ($p_{\mathrm{nl}}=1$) [30].
@@
The lack of consensus is largely due to a dearth of numerical and experimental guidance, which is missing because it is extremely challenging to accurately measure the nonlinear phase. At the lower values of Re accessible to numerical simulations and laboratory experiments, the nonlinear phase is hard to identify, due in part to the strong fluctuations of the integral quantities that can easily mask underlying trends. Our goal in this Letter is to directly measure $\alpha_{\mathrm{nl}}$ and polynomial growth order across a broad range of plasma parameters, allowing for the first time to confront theoretical predictions with highly-detailed measurements from simulations.
@@ II.1 Numerical Simulations
Figure 1: Volume-averaged magnetic energy $E_{\mathrm{mag}}$ normalized to the saturated value, $E_{\mathrm{mag},\mathrm{sat}}$ versus time for the subset of our SSD simulations with $\mbox{Re}=1500$; color shows the value of $\mathcal{M}$ for each run; the top panel shows the energy in logarithm, and the bottom, linear. We replicate each simulation configuration at least five times with different random seeds for the turbulent forcing. We plot median and 16th to 84th percentile range across these statistical realizations as the solid line and shaded band. The main axis shows time in units of outer-scale turbulent turnovers, $t_{0}$, and the inset axis repeats the bottom panel with time rescaled to the sound crossing time, $t_{\mathrm{sc}}$.
@@
We used a modified version of the flash code [39, 40, 41, 36] to run a series of direct numerical simulations that solve the compressible, non-ideal MHD equations for an isothermal plasma in a three-dimensional periodic box [24]; the full set of equations we solve is provided in End Matter. We work in dimensionless units where the box size $L$, mean density $\rho_{0}$, $c_{s}$, and mean thermal energy $\rho_{0}c_{s}^{2}$ are unity. The natural unit of time for this system is the sound-crossing time, $t_{\mathrm{sc}}=L/c_{\mathrm{s}}=1$, and the turbulent velocity $u_{0}$ directly sets the $\mathcal{M}=u_{0}/c_{s}$. By varying $u_{0}$, we control the speed of the flow relative to $t_{\mathrm{sc}}$, and in turn can explore subsonic ($t_{\mathrm{sc}}<t_{0}$) and supersonic ($t_{0}<t_{\mathrm{sc}}$) regimes. We utilize the turbgen [42] forcing-function to drive purely-solenoidal ($\bm{\nabla}\cdot\bm{f}=0$) turbulence with momentum source term $\bm{f}$, following the standard protocol outlined in Kriel et al. [23], allowing us to set $\mathcal{M}\in[5\times 10^{-2},5]$ (see End Matters for more details).
By varying the (constant in space and time) kinematic viscosity coefficient, $\nu$, we explore $\mbox{Re}\in[10^{3},5\times 10^{3}]$. Throughout this study, we set $\mbox{Pm}\equiv\nu/\eta=\mbox{Rm}/\mbox{Re}=1$, where $\eta$ is the magnetic resistivity coefficient and Rm is the magnetic Reynolds number. Because $\mbox{Re}=\mbox{Rm}\gtrsim\mbox{Rm}_{\mathrm{crit}}\approx 100$, all simulations are above the critical Rm to undergo all phases of the SSD [43, 38, 44]. In total we consider $12$ distinct (but over many different statistical realizations of $\bm{f}$, detailed in the next paragraph) combinations of $\mathcal{M}$ and Re – see End Matter for a complete list of simulations.
@@
In all our simulations we initialize a weak seed magnetic field $E_{\mathrm{mag},0}\equiv E_{\mathrm{mag}}(t=0)=10^{-10}E_{\mathrm{kin}}$, where $E_{\mathrm{kin}}$ is the kinetic energy once turbulence is fully established in the kinematic phase. We evolve each simulation instance long enough to unambiguously identify the onset and transition out of the nonlinear phase (see next section). To reduce the impact of fluctuations during the nonlinear phase on our fit parameters, we repeat each configuration at least five times with different random seeds for $\bm{f}$. We perform runs at three resolutions, $N_{\mathrm{res}}\in\{288^{3},576^{3},1152^{3}\}$. In total we run $89$ independent simulations. We plot examples of the time evolution for the integral magnetic field energy for a small selection ($\mbox{Re}=1500$) of these simulations in Fig. 1, where the classical SSD phases are clear to see.
@@ II.2 Fitting Procedure
To constrain growth timescales during the nonlinear phase, we begin by binning the raw time series of volume-averaged magnetic energy density $E_{\mathrm{mag}}$, for each of our simulations, into intervals of $t_{0}$. Within each time bin, we record the average and standard deviation of both $E_{\mathrm{mag}}$ and $\ln(E_{\mathrm{mag}})$. This yields, for each simulation $j$, a dataset $\bm{\mathcal{D}}_{j}\equiv(t_{i},\mu_{i},\mu_{i}^{(\ln)},\sigma_{i},\sigma_{i}^{(\ln)})$, where the data points $i$ are (approximately) uncorrelated, given that the turbulence correlation time is $t_{0}$.
Using a hierarchical two–stage strategy, we then employ a Bayesian approach to fit each $\bm{\mathcal{D}}_{j}$ to models for the time evolution of magnetic energy. In the first stage we only constrain the exponential and saturated behavior, and in the second we fit the full three–phase model with priors informed by the first stage. This ensures that the more complex nonlinear-phase dynamics remain stable and well-constrained, especially in our supersonic simulations, where large fluctuations can make the nonlinear phase difficult to identify. Anchoring in this way also improves the sampling efficiency.
The first stage of our model uses
@@
with $\bm{\theta}_{1}=(E_{0},\gamma_{\mathrm{exp}},t_{\mathrm{sat}})$, and the second stage uses
@@
where $\bm{\theta}_{2}=(E_{0},E_{\mathrm{sat}},\gamma_{\mathrm{exp}},t_{\mathrm{nl}},t_{\mathrm{sat}},p_{\mathrm{nl}})$. Note that $\alpha_{\mathrm{nl}}$ is not a fit parameter in the second stage, because it is implicitly set by the requirement that $f_{2}$ must remain continuous at $t=t_{\mathrm{sat}}$.
For our fits in both stages, we assume that the simulations represent Gaussian fluctuations around a mean trend, so for any given time series of simulation data $\bm{\mathcal{D}}_{j}$, the log likelihood function is
@@
where in stage one ($s=1$) we fit $(d_{i},e_{i})=(\mu_{i}^{(\ln)},\sigma_{i}^{(\ln)})$, and in stage two ($s=2$) we fit $(d_{i},e_{i})=(\mu_{i},\sigma_{i})$. This is because, as is apparent from Fig. 1, it is much easier to clearly identify the kinematic phase in logarithmic space, and the nonlinear phase in linear space.
We fit using a Markov Chain Monte Carlo method [45]; see End Matter for details on our sampling parameters. In the first stage we adopt uniform priors on $\gamma_{\mathrm{exp}}\in(0,10]$ and $t_{\mathrm{sat}}\in(0,t_{\mathrm{end}})$, where $t_{\mathrm{end}}$ is the final time bin, and a log-uniform prior on $E_{0}\in[10^{-30},10^{-5}]$ (due to its wide dynamic range). In the second stage, the priors for $E_{0}$, $\gamma_{\mathrm{exp}}$ and $E_{\mathrm{sat}}$ are taken to be the posteriors from stage one. We assign a uniform prior $p_{\mathrm{nl}}\in[1,2]$, consistent with existing theories, while, for $t_{\mathrm{nl}}$ and $t_{\mathrm{sat}}$ we use priors that are uniform on the intervals $0,t_{\mathrm{nl},\mathrm{max}}]$ and $t_{\mathrm{nl}},t_{\mathrm{sat},\mathrm{max}}]$, where $t_{\mathrm{nl},\mathrm{max}}$ is the time bin $t_{i}$ at which $dE_{\mathrm{mag}}/dt$ is its maximum, and $t_{\mathrm{sat},\mathrm{max}}$ is the earliest time $t_{i}$ for which $dE_{\mathrm{mag}}/dt<0$.
After fitting each individual simulation, we aggregate posteriors across runs with the same resolution and plasma parameters. This is the key step in our procedure, where we average out the inevitably large fluctuations in any single realization. Fitting each simulation separately preserves the phase structure of each realization, while still averaging over ensemble variability. This is crucial, because identical realizations may enter the nonlinear and saturated phases at slightly different times relative to one-another. Accurately constraining these transition times is essential for robustly measuring the nonlinear growth dynamics, which would be averaged out if all simulations were fit simultaneously. All the results we present in the remainder of this study are derived from these aggregated samples.

### v2

# The universal growth of magnetic energy during the nonlinear phase of subsonic and supersonic small-scale dynamos

@@ Abstract
Small-scale dynamos (SSDs) amplify magnetic fields in turbulent plasmas. Theory predicts nonlinear magnetic energy growth $E_{\mathrm{mag}}\propto t^{p_{\mathrm{nl}}}$, but this scaling has not been tested across flow regimes. Using a large ensemble of SSD simulations spanning subsonic to supersonic turbulence, we measure linear growth ($p_{\mathrm{nl}}=1$) in subsonic flows and quadratic growth ($p_{\mathrm{nl}}=2$) in supersonic flows. In all cases, the nonlinear dynamo converts a nearly constant fraction $\sim 1/100$ of the turbulent kinetic energy flux into magnetic energy, and the nonlinear phase has a characteristic duration $\Delta t\approx 20\,t_{0}$, where $t_{0}$ is the outer-scale turnover time. By isolating the onset of magnetic backreaction in SSDs, our statistical ensemble approach identifies a robust efficiency and duration for the nonlinear SSD that can be used to interpret more complex astrophysical and laboratory plasmas.
@@ I Introduction
Small-scale dynamo (SSD) action describes the process by which motions in a plasma with dynamically weak magnetic fields amplify and then maintain the magnetic energy density, $E_{\mathrm{mag}}$, to levels comparable to the turbulent kinetic energy density, $E_{\mathrm{kin}}$ [11, 35, 34]. This process is ubiquitous across astrophysical, geophysical, and laboratory environments, both those where the plasma supports turbulent subsonic motions (turbulent sonic Mach number $\mathcal{M}=u_{0}/c_{s}<1$, where $u_{0}$ is the rest-frame root-mean-squared velocity and $c_{s}$ the sound speed) and where they are supersonic ($\mathcal{M}>1$). The fields generated in this process magnetize the plasma between galaxies [36, 50], provide pressure support against collapse in galaxy mergers [55], enable rapid spin-down of stellar merger remnants [33, 37], and modify cosmic ray propagation via curvature acceleration in the interstellar medium (ISM) [24, 29]. Direct evidence for SSD action comes both from in situ observations in the Earth’s subsonic magnetosheath [53] and from supersonic laboratory laser experiments [31, 51, 6, 7].
@@
SSDs evolve through a number of distinct phases [32, 52, 28, 39, 40, 30, 35]. At first the magnetic field is too weak to exert significant forces on the plasma and thus the velocity field is kinematic, which leads to exponential amplification of the magnetic energy density, $E_{\mathrm{mag}}\propto\exp(\gamma_{\mathrm{exp}}t)$. Here $\gamma_{\mathrm{exp}}$ is the kinematic growth rate (with units of inverse time), and the dimensionless product $\gamma_{\mathrm{exp}}t_{0}$ measures how efficiently kinetic energy is converted into magnetic energy over one outer-scale eddy turnover time $t_{0}=\ell_{0}/u_{0}$, where $u_{0}$ is the flow velocity on the outer scale $\ell_{0}$ [32, 28]. For magnetic Prandtl numbers $\mbox{Pm}\gtrsim 1$, the dimensionless growth rate is predicted to scale as $\gamma_{\mathrm{exp}}t_{0}\propto\mbox{Re}^{1/2}$ for $\mathcal{M}<1$ and $\gamma_{\mathrm{exp}}t_{0}\propto\mbox{Re}^{1/3}$ for $\mathcal{M}>1$, where $\mbox{Re}\sim u_{0}\ell_{0}/\nu$ is the hydrodynamic Reynolds number [43] and $\nu$ is the kinematic viscosity. Field amplification is driven by chaotic stretching [35], which occurs on an eddy turnover timescale $t_{\ell}\sim\ell/u_{\ell}$, where $u_{\ell}$ is the flow speed on size scale $\ell$, and thus is fastest on the viscous scale $\ell_{\nu}$, which, for Kolmogorov turbulence, is the scale on which eddies evolve on the shortest timescales [5, 9, 22, 26, 25, 2]. However, once the kinematic dynamo amplifies the field such that $E_{\mathrm{mag}}\approx E_{\mathrm{kin}}$ on scale $\ell_{\nu}$, the magnetic field begins to back react on the velocity, marking the end of the kinematic phase and the start of the nonlinear phase. During this phase the dominant stretching scale $\ell_{s}$ shifts to larger $\ell$ with longer $t_{\ell}$, driving a slower, secular dynamo. As the field grows, it suppresses stretching on ever-larger scales, until $\ell_{s}$ reaches a maximum scale. At this point the dynamo reaches the saturated phase, and becomes statistically steady, driven MHD turbulence [1].
@@
Both numerical simulations [40, 26, 25], and laboratory experiments [51, 6, 7] confirm that the evolution of $E_{\mathrm{mag}}$ during the kinematic phase agrees well with both the Kazanstev, Anderson & Kulsrud SSD theory [23, 28] as well as with numerical works in the $\mathcal{M}>1$ regime [14, 17, 25]. However, this well-understood phase likely ends almost immediately in real astrophysical plasmas. This is because the growth rate is $\gamma_{\mathrm{exp}}t_{0}\propto\mbox{Re}^{1/2}$, and many warm or cold astrophysical plasmas, at least in our Galaxy and its surrounding medium, often have $\mbox{Re}\sim 10^{3}-10^{10}$, thus making the kinematic phase persist for only a small fraction of the outer dynamical timescale. Hence, almost any observable astrophysical system where the field is growing will instead be in the nonlinear phase, which is a regime that has not yet been explored with the same level of numerical detail as the kinematic dynamo.
@@
While there is indication that the nonlinear phase yields secular growth for the magnetic energy [4, 12, 46, 21, 25], $E_{\mathrm{mag}}=\alpha_{\mathrm{nl}}t^{p_{\mathrm{nl}}}$, there are no simulations to date that have measured the efficiency $\alpha_{\mathrm{nl}}$ and growth order $p_{\mathrm{nl}}$ across a wide range of plasma Reynolds number Re and Mach number $\mathcal{M}$. For incompressible Kolmogorov-like turbulence, where $E_{\mathrm{kin}}(\ell)\propto\ell^{5/3}$, the prevailing expectation and hint from numerical simulations is that $p_{\mathrm{nl}}=1$ [28, 39, 12, 4, 56], and $\alpha_{\mathrm{nl}}\propto\varepsilon$, where $\varepsilon\sim\rho_{\ell}u_{\ell}^{3}/\ell$ is the turbulent energy flux through scale $\ell$, with $\rho_{\ell}$ the density associated with motions on that scale (which is constant in the incompressible limit). In contrast, for highly-compressible, shock-dominated turbulence, where $E_{\mathrm{kin}}(\ell)\propto\ell^{2}$ (as in Burgers-like turbulence [18, 16, 25, 2]), the outcome is debated: some models [41, 44] predict quadratic growth ($p_{\mathrm{nl}}=2$), while others suggest universal linear growth ($p_{\mathrm{nl}}=1$) [4].
@@
This lack of consensus about the nonlinear growth order $p_{\mathrm{nl}}$ and efficiency $\alpha_{\mathrm{nl}}$, particularly in highly-compressible turbulence, is largely due to a dearth of numerical and experimental guidance, which is missing in large part because it is extremely challenging to accurately measure the nonlinear phase. At the modest values of Re accessible to numerical simulations and laboratory experiments, the nonlinear phase is hard to identify, due in part to the strong fluctuations of integral quantities that can easily mask underlying trends, especially in supersonic flows. Our goal in this study, is to take on these challenges by focusing on plasmas with $\mbox{Pm}=1$, for which the nonlinear phase we resolve, can only be the backreaction stage that begins once $E_{\mathrm{mag}}\approx E_{\mathrm{kin}}$ on the viscous scale. This choice also maximizes the computational dynamic range over which the nonlinear backreaction can operate, because the viscous and resistive dissipation scales for any choice of Re or Rm coincide with one-another ($\ell_{\nu}\approx\ell_{\eta}$). Within this setting we directly measure $\alpha_{\mathrm{nl}}$ and $p_{\mathrm{nl}}$ across a broad range of plasma parameters (in Re and $\mathcal{M}$), allowing us, for the first time, to confront theoretical predictions for the backreaction during the nonlinear phase with highly-detailed measurements from simulations. At the same time, our results provide a well-controlled reference point for future studies of more realistic hot astrophysical plasmas with $\mbox{Pm}\gg 1$, where this backreaction stage may be preceded by an additional nonlinear, secular growth phase in the sub-viscous range (see the discussion in the Conclusion).
@@ II.1 Numerical Simulations
Figure 1: Time evolution of the volume-averaged magnetic energy $E_{\mathrm{mag}}$, normalized by the final saturated value $E_{\mathrm{mag},\mathrm{sat}}$, for a subset of SSD simulations with $\mbox{Re}=1500$. The top panel shows the energy in logarithmic scale, and the bottom panel (and inset) shows it in linear scale. Each simulation configuration is repeated five times with independent random seeds for the turbulent forcing; we plot the median trend across realizations as a solid line, with the 16th to 84th percentile range indicated by a shaded band. Colors indicate the average sonic Mach number $\mathcal{M}$ of each configuration. In the main axes, time is shown in units of the turbulent turnover time on the outer scale, $t_{0}$, and rescaled to the sound-crossing time, $t_{\mathrm{sc}}$, in the inset.
@@
We used a modified version of the flash code [20, 13, 54, 16] to run a series of direct numerical simulations that solve the compressible, non-ideal MHD equations for an isothermal plasma in a three-dimensional periodic box [25]; the full set of equations we solve is provided in End Matter. We work in dimensionless units where the box size $L$, mean density $\rho_{0}$, $c_{s}$, and mean thermal energy $\rho_{0}c_{s}^{2}$ are unity. The natural unit of time for this system is the sound-crossing time, $t_{\mathrm{sc}}=L/c_{\mathrm{s}}=1$, and the turbulent velocity $u_{0}$ directly sets the sonic Mach number, $\mathcal{M}=u_{0}/c_{s}$. By varying $u_{0}$, we control the speed of the flow relative to $t_{\mathrm{sc}}$, and in turn can explore subsonic ($t_{\mathrm{sc}}<t_{0}$) and supersonic ($t_{0}<t_{\mathrm{sc}}$) regimes. We use the turbgen [15] forcing-module to drive purely-solenoidal ($\bm{\nabla}\cdot\bm{f}=0$) turbulence with momentum source term $\bm{f}$, following the standard protocol outlined in Kriel et al. [26], allowing us to set $\mathcal{M}\in[5\times 10^{-2},5]$ (see End Matter for more details).
By varying the (constant in space and time) kinematic viscosity coefficient, $\nu$, across different simulation configurations, we explore a range of hydrodynamic Reynolds numbers, $\mbox{Re}\in[10^{3},5\times 10^{3}]$, where we set the magnetic Prandtl number $\mbox{Pm}\equiv\nu/\eta=\mbox{Rm}/\mbox{Re}=1$; $\eta$ is the magnetic resistivity coefficient and Rm is the magnetic Reynolds number. We quote Re and Rm using $u_{0}$, which is the rest-frame root-mean-squared velocity measured in the statistically steady kinematic phase. For convenience, we use these kinematic values to label simulations and compare across the parameter space. During the nonlinear and saturated phases, magnetic backreaction slightly suppresses the velocity amplitude, so the instantaneous effective values of Re and Rm may be smaller than their nominal kinematic values. This effect, however, is minor for $\mbox{Pm}=1$, becoming more prominent for $\mbox{Pm}\gg 1$ [47, 48]. Because $\mbox{Re}=\mbox{Rm}\gtrsim\mbox{Rm}_{\mathrm{crit}}\approx 100$, all simulations are above the critical Rm required to undergo all phases of the SSD [38, 44, 45]. In total we consider $12$ distinct (but over many different statistical realizations of $\bm{f}$, detailed in the next paragraph) combinations of $\mathcal{M}$ and Re – see End Matter for a complete list of simulations. We also ran pilot simulations at lower Re, but found that the volume-integrated kinetic and magnetic energies had relative fluctuations that were too large for our analysis. When the viscous and resistive dissipation scales $\ell_{\nu}$ and $\ell_{\eta}$ lie close to the forcing scale $\ell_{0}$, the cascade is truncated and fluctuations remain confined to large scales. The effective number of independent eddies contributing to volume averages, $N_{\rm eddies}\sim(L/\ell_{\nu})^{3}$, is then small, so the fractional fluctuations $\propto N_{\rm eddies}^{-1/2}$ remain large in the nonlinear phase. For this reason we adopt moderately turbulent flows with $\mbox{Re}\gtrsim 10^{3}$ as our practical lower bound. Finally, we caution readers that there are two different conventions in use in the literature to define the Reynolds number; our definition, $\mbox{Re}\sim u_{0}\ell_{0}/\nu$, is a factor of $2\pi$ larger than that used by some authors [e.g.,  10, 21], so in the convention used by these authors, our simulations span $\mbox{Re}\in[1.5\times 10^{2},8\times 10^{2}]$.
@@
In all our simulations we initialize a weak seed magnetic field $E_{\mathrm{mag},0}\equiv E_{\mathrm{mag}}(t=0)=10^{-10}E_{\mathrm{kin}}$, where $E_{\mathrm{kin}}$ is the kinetic energy once turbulence is fully established in the kinematic phase. We evolve each simulation instance long enough to unambiguously identify the onset and transition out of the nonlinear phase (see next section). To reduce the impact of fluctuations during the nonlinear phase on our fit parameters, we repeat each configuration at least five times with different random seeds for $\bm{f}$. We perform runs at three resolutions, $N_{\mathrm{res}}\in\{288^{3},576^{3},1152^{3}\}$, and run a total of $89$ independent simulation instances. We plot the time evolution for the integral magnetic field energy of a representative subset of our simulations ($\mathcal{M}$-varied with $\mbox{Re}=1500$) in Fig. 1, where the classical SSD phases are clear to see. The main panels use $t_{0}$ to emphasize the evolution in outer-scale turnover times, while the inset re-normalizes the same curves in units of the sound-crossing time $t_{\mathrm{sc}}$, highlighting how the ordering of the nonlinear phase onset changes when measured against $t_{\mathrm{sc}}$ rather than $t_{0}$.
@@ II.2 Fitting Procedure
To constrain growth timescales during the nonlinear phase, we begin by binning the raw time series of volume-averaged magnetic energy density $E_{\mathrm{mag}}$, for each of our simulations, into intervals of $t_{0}$. Within each time bin $t_{i}$, we compute the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ of both $E_{\mathrm{mag}}$ and $\ln(E_{\mathrm{mag}})$ in bin $i$; $\mu_{i}^{(\ln)}$ and $\sigma_{i}^{(\ln)}$ indicate the mean and standard deviation of $\ln(E_{\mathrm{mag}})$ in bin $i$. This yields, for each simulation $j$, a dataset $\bm{\mathcal{D}}_{j}\equiv(t_{i},\mu_{i},\mu_{i}^{(\ln)},\sigma_{i},\sigma_{i}^{(\ln)})$, where data points are (approximately) uncorrelated, given that the turbulence correlation time is $t_{0}$.
Using a hierarchical two-stage strategy, we then employ a Bayesian approach to fit each $\bm{\mathcal{D}}_{j}$ to models for the time evolution of magnetic energy. In the first stage we only constrain the exponential and saturated behavior, and in the second stage we fit the full three-phase model with priors informed by the first. This ensures that the more complex nonlinear-phase dynamics remain stable and well-constrained, especially in our supersonic simulations, where large fluctuations can make the nonlinear phase difficult to identify. Anchoring in this way also improves the sampling efficiency.
Our first stage models only the kinematic and saturated phases, approximating the transition between these phases as instantaneous,
@@
with $\bm{\theta}_{1}=(E_{0},\gamma_{\mathrm{exp}},t_{\mathrm{sat}})$, and the second stage models all dynamo phases self-consistently as
@@
where $\bm{\theta}_{2}=(E_{0},E_{\mathrm{sat}},\gamma_{\mathrm{exp}},t_{\mathrm{nl}},t_{\mathrm{sat}},p_{\mathrm{nl}})$. Here, $\alpha_{\mathrm{nl}}$ is not a fit parameter in the second stage, because it is implicitly set by the requirement that $f_{2}$ must remain continuous at $t=t_{\mathrm{sat}}$.
For our fits in both stages, we assume that each data point $d_{i}$ in the time series is an independent sample from a Gaussian distribution around a mean trend, so for any given time series of simulation data $\bm{\mathcal{D}}_{j}$, the log likelihood function is
@@
where we fit $(d_{i},e_{i})=(\mu_{i}^{(\ln)},\sigma_{i}^{(\ln)})$ in stage one ($s=1$), and we fit $(d_{i},e_{i})=(\mu_{i},\sigma_{i})$ in stage two ($s=2$). This is because, as is apparent from Fig. 1, it is much easier to identify the kinematic (exponential growing) phase in logarithmic space, and the nonlinear (backreaction) phase in linear space.
We fit using a Markov Chain Monte Carlo method [19]; see End Matter for details on our sampling parameters. In the first stage we adopt uniform (unbiased) priors on $\gamma_{\mathrm{exp}}\in(0,10]$ and $t_{\mathrm{sat}}\in(0,t_{\mathrm{end}})$, where $t_{\mathrm{end}}$ is the final time bin, and a log-uniform prior on $E_{0}\in[10^{-30},10^{-5}]$ (due to its wide dynamic range). In the second stage, the priors for $E_{0}$, $\gamma_{\mathrm{exp}}$ and $E_{\mathrm{sat}}$ are taken to be the posteriors from stage one. We assign a uniform prior $p_{\mathrm{nl}}\in[1,2]$, consistent with existing theories, while, for $t_{\mathrm{nl}}$ and $t_{\mathrm{sat}}$ we use priors that are uniform on the intervals $0,t_{\mathrm{nl},\mathrm{max}}]$ and $t_{\mathrm{nl}},t_{\mathrm{sat},\mathrm{max}}]$, where $t_{\mathrm{nl},\mathrm{max}}$ is the time bin where $dE_{\mathrm{mag}}/dt$ is its maximum, and $t_{\mathrm{sat},\mathrm{max}}$ is the earliest time $t_{i}$ for which $dE_{\mathrm{mag}}/dt<0$.
After fitting each individual simulation, we aggregate posteriors across runs with the same resolution and plasma parameters. This is the key step in our procedure, where we average out the inevitably large fluctuations in any single realization. For a given configuration, all realizations are statistically equivalent and differ only in their random driving, so we combine their posterior samples with equal weighting. Fitting each simulation separately preserves the phase structure of each realization, while still averaging over ensemble variability. This is crucial, because identical realizations may enter the nonlinear and saturated phases at slightly different times relative to one another. Accurately constraining these transition times is essential for robustly measuring the nonlinear growth dynamics, which would be averaged out if all instances of a configuration were fit simultaneously. All the results we present in the remainder of this study are derived from these aggregated samples; percentiles of these derived quantities are reported in Table 1, and the raw data, along with our implementation of the routines discussed here, are publicly available online [27].